If 1/3 of an office staff is on vacation and 1/4 of those not on vacation are out sick, what fraction of the staff is in the office?
Answer by fcabanski(1391)

If 1/3 of the staff is on vacation, then 3/3 -1/3 = 2/3 remain.


1/4 of (of means times) the remaining 2/3 is sick. 1/4 * 2.3 = 2/12 are sick.


That leaves 2/3 - 2/12 = 8/12-2/12 = 6/12 = 1/2 of the staff in the office.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)